# Machine learning of the Ising model on a spherical Fibonacci lattice

### v1

@@ Abstract
We investigate the Ising model confined to a spherical surface, focusing on its implementation using a Fibonacci lattice. The challenge lies in uniformly covering the spherical surface to enable reliable comparisons with planar models. Monte Carlo simulations and graph convolutional networks(GCNs) are employed to analyze spin configurations at varying temperatures and to identify phase transition temperatures. Although the spherical Fibonacci lattice is sufficiently uniform, there are still some irregular sites, which introduce interesting effects. In the ferromagnetic case, sites with fewer neighbors are more likely to undergo spin flips at low temperatures; however, this is not necessarily true at high temperatures, which could explain why the phase transition temperature is higher compared to the planar Ising model. In the antiferromagnetic case, the presence of irregular sites results in the total energy of the system at zero temperature not being the lowest. Phase transition temperatures are estimated using specific heat analysis and GCNs, revealing $T_{C}$ values for both ferromagnetic and antiferromagnetic cases. The study underscores the significance of the Fibonacci lattice’s geometric properties in understanding spin interactions in microgravity environments.
@@ I Introduction
The first challenge of this problem is to cover the spherical surface as uniformly as possible. Only in this way can we reliably compare the results with those of the planar square lattice. This requirement excludes the traditional latitude-longitude lattice, as the density of sites near the poles is significantly higher than anywhere else. The most suitable candidate is the Fibonacci lattice, which is essentially the most uniform spherical lattice. We have previously applied it to study the spherical XY model [25], yielding intriguing results about how vortex distribution is significantly influenced by spherical topology. It is important to note that the Fibonacci lattice is not perfectly uniform, which complicates the analytical study of the Ising model on the Fibonacci sphere. However, by utilizing Monte Carlo(MC) simulations and machine learning techniques, we can determine the spin configurations at various temperatures and identify the phase-transition temperature of the spherical Ising model. The second challenge is that the Fibonacci lattice, being an irregular lattice, lacks an image-like structure. Therefore, powerful tools such as convolutional neural networks (CNNs) cannot be applied directly to the Fibonacci lattice. Instead, we use graph convolutional neural networks(GCNs) to handle irregular lattices, a method that we have previously applied to study the XY model on a spherical Fibonacci lattice[25]. GCNs are effective in classifying different spin phases and successfully predicting the phase-transition temperature. Recently, a study focused on the ferromagnetic Ising model on a Fibonacci-triangulated sphere found that the model exhibits a critical temperature slightly lower than that of a planar triangular lattice[26]. In this paper, we will concentrate on the mostly quadrangulated Fibonacci lattice for both ferromagnetic and antiferromagnetic Ising models, where neighbor interactions are determined by a cutoff radius $r_{c}$.
The rest of this paper is organized as follows: In SectionII, we construct the spherical Fibonacci lattice points, ensuring that the number of lattice sites with four nearest neighbors is maximized by selecting an appropriate spin-interaction cutoff radius $r_{c}$. SectionsIII and IV present simulations of the spherical ferromagnetic and antiferromagnetic Ising models, respectively, using MC algorithms and GCNs. These sections include graphs illustrating spin configurations, energy and specific heat at various temperatures, as well as the phase transition temperatures for both models. Finally, SectionV concludes the paper.
@@ II Spherical Ising Model
$H_{\{s_{i}\}}=-J\sum_{\langle i,j\rangle}s_{i}s_{j}-h\sum_{i}s_{i},$ (1)
where $s_{i}$ represents the spin at site $i$, $J$ is the interaction strength and $h$ is the external magnetic field. The first summation runs over all pairs of adjacent spins, with each pair being counted only once. In this paper, we will focus on the case with $h=0$ for convenience. Ising models can be classified according to the sign of $J$: If $J>0$, the interaction is called ferromagnetic; if $J<0$, it is antiferromagnetic. For irregular lattices, the distance between adjacent spins may not be a constant. Therefore, it is crucial to carefully define the ”nearest-neighbor sites”, which will be discussed in detail later.
@@
$z_{i}$ $=R(\frac{2i-1}{N}-1),$ (2)
where the notation $i=1,2,\cdots,N$ denotes the index of a lattice point, $R$ represents the radius of the sphere, and $\phi=\frac{\sqrt{5}-1}{2}$ is the golden ratio. Fig.1 compares the traditional latitude-longitude lattice with the Fibonacci lattice. Evidently, the latter is considerably more uniform than the former. Next, to compare with the properties of the Ising model on the planar square lattice, we quadrangulate the spherical Fibonacci lattice as much as possible. Equivalent, this means that most spins interact with four nearest neighbors. To achieve this, we set a critical radius $r_{c}$. If the distance between two neighboring spins is less than $r_{c}$, an interaction is considered to exist between them. To maximize the number of lattice sites with four nearest neighbors in a system with $N=1000$ lattice sites, we determined the optimal nearest-neighbor radius through exploratory analysis, setting it to $r_{c}=0.1298R$. In this situation, there are $850$ spins with $4$ neighbors, $76$ spins with $3$ neighbors, and $74$ spins with $5$ neighbors. Overall, the $1000$ spins have a total of $3998$ neighbors, which is very close to the square lattice of $N=1000$. We connect all ”nearest-neighbor sites” and present a two perspectives of the quadrangulation of a $N=1000$ Fibonacci lattice from two different directions in Fig.2. Most sites appear ”regular” except a few having $3$ or $5$ neighbors. In practical experiments, implementing a Fibonacci lattice is quite straightforward. We simply need to distribute the lattice points as evenly as possible on a spherical surface. This arrangement will naturally approximate a Fibonacci lattice in an appropriate coordinate system, as the Fibonacci lattice is fundamentally the most uniform lattice on a sphere.
@@ III Ferromagnetic Ising model on a spherical Fibonacci lattice
At low temperatures, it is evident that all sites have the same spin orientation. As the temperature starts to increase, thermal fluctuations may cause spins at certain sites to reverse direction. However, it is not until the temperature reaches around $T/J\sim 1.9$ that the number of reversed spins becomes fairly considerable. In the top panel of Fig.3, we plot the stable spin configuration at a relatively low temperature of $T/J=2.0$, where up and down spins are represented by white and black pixels, respectively. It is observed that some spins are flipped(black points), while the majority of the regions remain predominantly occupied by up spins(white points). Clearly, the system is in the ordered phase. At a very high temperature of $T/J=8.0$, a significant number of spins are flipped due to thermal fluctuations. As a result, the numbers of up and down spin sites become roughly comparable. The system is now in the disordered phase, and the corresponding spin configuration is shown in the bottom panel of Fig.3.
@@
Figure 4: (Left panel) A perspective of the Fibonacci lattice at $T/J=2.0$, in which the spins at $6$ sites with three neighbours change direction. (Right panel) Statistical spin-flip probabilities for spin-flip as a function of temperature.
@@
Note that the Fibonacci lattice is inhomogeneous, with varying probabilities of spin flips at different sites, which is crucial for understanding stable spin patterns. For instance, a site with three neighbors incurs an energy cost of $6J$ to flip its spin, while a site with five neighbors incurs a cost of $10J$. Consequently, the probability of a spin-flip is relatively higher for the site with three neighbors. In Fig.4, we present a view of the Fibonacci lattice from a particular direction at $T/J=2.0$. It is evident that among the sites where the spin is flipped, at least $6$ out of $76$ sites have three neighbors. In contrast, only $11$ out of $850$ sites with four neighbors exhibit spin flips. This indicates that the proportion of spin flips is relatively significant for sites with three neighbors. We also statistically calculate the average the spin-flip probabilities $P_{3}$, $P_{4}$ and $P_{5}$ for sites with $3$, $4$ and $5$ neighbours, respectively, across different stable spin configurations at various temperatures. The results are illustrated in Fig.4 by black squares representing $P_{3}$, red disks representing $P_{4}$, and blue triangles representing $P_{5}$. Specifically, we perform ten simulations at each temperature. For each simulation, we count the number of spin flips occurring at the sites with $3$, $4$, and $5$ neighbours in a stable spin configuration. We then compute the ratio of these flips to the total number of sites for each type to obtain the flip ratio. Finally, we average the results from the ten simulations at each temperature. At low temperatures, $P_{3}>P_{4}>P_{5}$, reflecting that sites with fewer neighbors are more prone to spin flips. At the temperature increases, thermal fluctuations become more significant relative to the energy cost of spin flips, leading to a disruption in the ordering of $P_{3}$, $P_{4}$ and $P_{5}$.
@@
Figure 5: (Left panel) Total energy as a function of temperature for the spherical ferromagnetic Ising model with $N=1000$. (Right panel) Specific heat as a function of temperature for the same model.
@@
In the end, we aim to determine the phase transition temperature $T_{c}$ of the spherical Fibonacci Ising model. To achieve this, we employ two different methods. The first method is straightforward but relatively approximate: We estimate $T_{c}$ by analyzing the behavior of the specific heat $C_{V}$. The second method offers greater accuracy and involves using graph convolutional networks(GCN) to refine our estimation of $T_{c}$.
@@
Given a stable spin configuration, the total energy $E$ of the system is calculated via Eq.(1). Similarly, we average the total energy over $10$ stable spin configuration at each temperature, with a temperature interval $0.01J$. The numerical results are presented in the left panel of Fig.5. To estimate the specific heat, we use the spine functions to fit the $E$ versus $T$ curve, resulting in a function $E(t)$. We then compute the derivative of this function to estimate $C_{V}$ and plot $C_{V}$ vs $T$ in the right panel of Fig.5. As shown in the figure, $C_{V}$ exhibits a singular behavior at $T/J=2.29$, which can be considered as an estimate of $T_{c}$. Interestingly, this value is very close to the critical temperature of the two-dimensional Ising model on a square lattice, which is $T^{\Box}_{c}/J\simeq 2.269$[30].
@@
The second approach employs machine learning, which has proven to be a powerful tool in processing big data[31, 32, 33, 34, 35, 36, 37, 38]. Given that the Fibonacci lattice is not homogeneous and does not resemble a typical image-like structure, traditional CNNs can not be applied to this model[39, 40]. Instead, the GCN will be utilized, as it is capable of capturing spatial structural features by leveraging the connectivity relationships between nodes[41]. This makes GCNs particularly well-suited for processing topologically structured data like the Fibonacci lattice. Details of GCN can be found in AppendixA.
We utilize the results from MC simulations as the training samples at both low and very high temperature regimes. The GCN can be subsequently used to identify ordered and disordered phases, as well as phase transitions. To verify the accuracy of the GCN simulations, we investigate two-dimensional square lattices with sizes of $20\times 20$ and $40\times 40$. The algorithm generates classification confidence values for the ordered phases($p_{\text{o}}$) and the disordered phases ($p_{\text{d}}$), as illustrated in Fig.(6) (with the black solid line representing $p_{\text{o}}$ and the red dashed line representing $p_{\text{d}}$). As the temperature increases, the black solid line shows a gradual decline from $1$, indicating a decreasing similarity between the simulation results and the ordered phase. In contrast, the red dashed line gradually increases from 00, indicating an enhancement in the similarity between the simulation results and the disordered phase. The temperature at which $p_{\text{o}}=p_{\text{d}}$ signifies the phase transition temperature $T_{c}$. Through the simulation and analyses of different lattice scales, it is found that as the number of lattice sites increases, the phase transition temperature gradually approaches the theoretical value of $T_{c}/J\simeq 2.269$. Specifically, the values obtained are $T_{c}/J=2.51$ for the $20\times 20$ lattice, and $T_{c}/J=2.36$ for the $40\times 40$ lattice. The latter demonstrates a strong concordance with the predicted critical temperature. This indicates that increasing the lattice size can enhance the accuracy of phase transition temperature estimations. Moreover, the GCN effectively captures the phase transition behavior and critical properties.
After validating the accuracy of GCN, we can subsequently use it to calculate the phase transition temperature of the spherical Ising model. In numerical calculation, we select 791 temperature nodes within the range of $[0.1,8]J$, with an interval of $0.01J$, and perform $100$ simulations for each temperature node. The results, shown in Fig.7, indicate that the critical temperature is determined to be $T_{c}/J=2.44$. This value is comparatively higher than that obtained from the analysis of the specific heat $C_{V}$. Due to the inhomogeneity of the Fibonacci lattice, a complete analytical analysis is not feasible. We believe a possible reason for this discrepancy is the presence of sites with $3$ or $5$ neighbors, which affects the spin dynamics within the system. At low temperatures, a site with a greater number of neighbors is less likely to experience a spin flip, as verified in Fig.4. However, as the temperature increases, the situation changes. Fig.4 shows that $P_{4}>P_{3}>P_{5}$ in the rangle $2.4\lesssim T/J\lesssim 2.5$, which indicating that the spins on sites with $3$ or $5$ neighbors are more difficult to flip. This, to some extent, increases the ’stability’ of the ordered phase, which elevating the phase transition temperature.
@@
Figure 6: Classification confidences $p_{o}$ and $p_{d}$ versus temperature for the 2D Ising model on a $20\times 20$ square lattice(Left panel) and $40\times 40$ square lattice(Right panel).

### v2

@@ Abstract
We investigate the Ising model on a spherical surface, utilizing a Fibonacci lattice to approximate uniform coverage. This setup poses challenges in achieving consistent lattice distribution across the sphere for comparison with planar models. We employ Monte Carlo simulations and graph convolutional networks (GCNs) to study spin configurations across a range of temperatures and to determine phase transition temperatures. The Fibonacci lattice, despite its uniformity, contains irregular sites that influence spin behavior. In the ferromagnetic case, sites with fewer neighbors exhibit a higher tendency for spin flips at low temperatures, though this effect weakens as temperature increases, leading to a higher phase transition temperature than in the planar Ising model. In the antiferromagnetic case, lattice irregularities prevent the total energy from reaching its minimum at zero temperature, highlighting the role of curvature and connectivity in shaping interactions. Phase transition temperatures are derived through specific heat, magnetic susceptibility analysis and GCN predictions, yielding $T_{c}$ values for both ferromagnetic and antiferromagnetic scenarios. This work emphasizes the impact of the Fibonacci lattice’s geometric properties-namely curvature and connectivity-on spin interactions in non-planar systems, with relevance to microgravity environments.
@@ I Introduction
The first challenge of this problem is to cover the spherical surface as uniformly as possible. Only in this way can we reliably compare the results with those of the planar square lattice. This requirement excludes the traditional latitude-longitude lattice, as the density of sites near the poles is significantly higher than anywhere else. The most suitable candidate is the Fibonacci lattice, which is essentially the most uniform spherical lattice. We have previously applied it to study the spherical XY model [25], yielding intriguing results about how vortex distribution is significantly influenced by spherical topology. It is important to note that the Fibonacci lattice is not perfectly uniform, which complicates the analytical study of the Ising model on the Fibonacci sphere. However, by utilizing Monte Carlo (MC) simulations and machine learning techniques, we can determine the spin configurations at various temperatures and identify the phase-transition temperature of the spherical Ising model.
The second challenge arises from the irregular nature of the Fibonacci lattice, which lacks an image-like structure. Machine learning techniques have been widely used to identify phase transitions in statistical physics models, such as the Ising model. Notably, Carrasquilla and Melko [26] demonstrated the effectiveness of supervised learning with convolutional neural networks (CNNs) in classifying ordered and disordered phases in the two-dimensional square-lattice Ising model, successfully determining the critical temperature from labeled configurations. However, this approach cannot be directly applied to the Fibonacci lattice due to its non-uniform geometry. To address this, we extend the paradigm of machine learning for phase transition detection to the irregular geometry of the spherical Fibonacci lattice. Specifically, we employ graph convolutional networks (GCNs), a natural generalization of CNNs for non-Euclidean structures, which we have previously applied to study the XY model on a spherical Fibonacci lattice [25].
Recently, a study focused on the ferromagnetic Ising model on a Fibonacci-triangulated sphere found that the model exhibits a critical temperature slightly lower than that of a planar triangular lattice [27]. In this paper, we will concentrate on the mostly quadrangulated Fibonacci lattice for both ferromagnetic and antiferromagnetic Ising models, where neighbor interactions are determined by a cutoff radius $r_{c}$. Through the application of specific heat analysis and GCNs, we determined the phase transition temperatures for both situations, highlighting the influence of geometric properties on spin interactions.
The rest of this paper is organized as follows: In Section II, we construct the spherical Fibonacci lattice points, ensuring that the number of lattice sites with four nearest neighbors is maximized by selecting an appropriate spin-interaction cutoff radius $r_{c}$. Sections III and IV present simulations of the spherical ferromagnetic and antiferromagnetic Ising models, respectively, using MC algorithms and GCNs. These sections include graphs illustrating spin configurations, energy, entropy and specific heat at various temperatures, as well as the phase transition temperatures for both models. Finally, Section V concludes the paper.
@@ II Spherical Ising Model
$H_{\left\{s_{i}\right\}}=-J\sum_{\langle i,j\rangle}s_{i}s_{j}-h\sum_{i}s_{i},$ (1)
where $s_{i}$ represents the spin at site $i$, $J$ is the interaction strength and $h$ is the external magnetic field. The first summation runs over all pairs of adjacent spins, with each pair being counted only once. In this paper, we will focus on the case with $h=0$ for convenience. Ising models can be classified according to the sign of $J$: If $J>0$, the interaction is called ferromagnetic; if $J<0$, it is antiferromagnetic. For irregular lattices, the distance between adjacent spins may not be a constant. Therefore, it is crucial to carefully define the “nearest-neighbor sites”, which will be discussed in detail later.
@@
$z_{i}$ $=R\left(\frac{2i-1}{N}-1\right),$ (2)
where the notation $i=1,2,\cdots,N$ denotes the index of a lattice point, $R$ represents the radius of the sphere, and $\phi=\frac{\sqrt{5}-1}{2}$ is the golden ratio. Fig.1 compares the traditional latitude-longitude lattice with the Fibonacci lattice. Evidently, the latter is considerably more uniform than the former. Next, to compare with the properties of the Ising model on the planar square lattice, we quadrangulate the spherical Fibonacci lattice as extensively as possible. In essence, this means that the majority of spins interact with four nearest neighbors. To achieve this, we set a critical radius $r_{c}$. If the distance between two neighboring spins is less than $r_{c}$, an interaction is considered to exist between them. Firstly, we consider a system with $N=1000$ lattice sites as an example. To maximize the number of lattice sites with four nearest neighbors, we perform an exploratory analysis to determine the optimal nearest-neighbor radius, setting it to $r_{c}=0.1298R$. Under these conditions, 850 spins have four neighbors, 76 spins have three neighbors, and 74 spins have five neighbors. In total, the 1000 spins collectively have 3998 neighboring connections, closely approximating the structure of a square lattice with $N=1000$. We connect all “nearest-neighbor sites” and present a two perspectives of the quadrangulation of a $N=1000$ Fibonacci lattice from two different directions in Fig.2. Most sites appear “regular” except a few having 3 or 5 neighbors. In practical experiments, implementing a Fibonacci lattice is quite straightforward. We simply need to distribute the lattice points as evenly as possible on a spherical surface. This arrangement will naturally approximate a Fibonacci lattice in an appropriate coordinate system, as the Fibonacci lattice is fundamentally the most uniform lattice on a sphere.
@@ III.1 Spin configurations
At low temperatures, it is evident that all sites have the same spin orientation. As the temperature starts to increase, thermal fluctuations may cause spins at certain sites to reverse direction. In the top panel of Fig.3, we plot the stable spin configuration at a relatively low temperature of $T/J=2.0$, where up and down spins are represented by white and black pixels, respectively. It is observed that some spins are flipped (black points), while the majority of the regions remain predominantly occupied by up spins (white points). Clearly, the system is in the ordered phase. At a very high temperature of $T/J=8.0$, a significant number of spins are flipped due to thermal fluctuations. As a result, the numbers of up and down spin sites become roughly comparable. The system is now in the disordered phase, and the corresponding spin configuration is shown in the bottom panel of Fig. 3.
@@
Figure 4: (Top panel) A perspective of the Fibonacci lattice at $T/J=2.0$, in which the spins at 6 sites with three neighbours change direction (Left). The ranking of the statistical spin-flip probabilities $P_{3}$, $P_{4}$ and $P_{5}$ at different temperatures (Right). (Bottom panel) Statistical spin-flip probabilities as functions of temperature.
@@
Note that the Fibonacci lattice is inhomogeneous, with varying probabilities of spin flips at different sites, which is crucial for understanding stable spin patterns. For instance, a site with three neighbors incurs an energy cost of $6J$ to flip its spin, while a site with five neighbors incurs a cost of $10J$. Consequently, the probability of a spin-flip is relatively higher for the site with three neighbors. In the left subpanel of the top row in Fig.4, we present a view of the Fibonacci lattice from a particular direction at $T/J=2.0$. It is evident that among the sites where the spin is flipped, at least 6 out of 76 sites have three neighbors. In contrast, only 11 out of 850 sites with four neighbors exhibit spin flips. This indicates that the proportion of spin flips is relatively significant for sites with three neighbors. We also statistically calculate the average spin-flip probabilities $P_{3}$, $P_{4}$ and $P_{5}$ for sites with 3, 4 and 5 neighbours, respectively, across different stable spin configurations at various temperatures. The results are illustrated in the bottom panel of Fig.4, with black squares representing $P_{3}$, red disks representing $P_{4}$, and blue triangles representing $P_{5}$. To better illustrate the temperature-dependent changes in the relative spin-flip probabilities for different neighboring sites, we display the ranking of $P_{3}$, $P_{4}$ and $P_{5}$ as a function of temperature in the right subpanel of the top row in Fig.4, with black, red and blue squares representing $P_{3}$, $P_{4}$, and $P_{5}$ respectively. The numbers on the left, ranging from 1 to 3, indicate the relative ranking of the values rather than their specific numerical values. Specifically, we perform ten simulations at each temperature. For each simulation, we count the number of spin flips occurring at the sites with 3, 4, and 5 neighbours in a stable spin configuration. We then compute the ratio of these flips to the total number of sites for each type to obtain the flip ratio. Finally, we average the results from the ten simulations at each temperature. At low temperatures, $P_{3}>P_{4}>P_{5}$, reflecting that sites with fewer neighbors are more prone to spin flips. As the temperature increases to $T/J\gtrsim 2.34$, thermal fluctuations become more significant relative to the energy cost of spin flips, leading to a disruption in the ordering of $P_{3}$, $P_{4}$ and $P_{5}$. The inset of the bottom panel displays that $P_{4}>P_{5}>P_{3}$ at $T/J=2.34$. This effect has a significant impact on the phase transition temperature, as will be demonstrated later.
@@
Figure 5: (Top panel) Total energy as a function of temperature for the spherical ferromagnetic Ising model with $N=1000$. (Middle panel) Pecific heat as a function of temperature, calculated using $C_{V}=\left(\frac{\partial E}{\partial T}\right)_{V}$. (Bottom panel) Specific heat as a function of temperature, calculated using Eq.(3).
III.2 Determining $T_{c}$ via the behaviors of the specific heat and the magnetic susceptibility
@@
Next, we aim to determine the phase transition temperature $T_{c}$ of the spherical Fibonacci Ising model. To achieve this, we employ three different approaches. The first one is straightforward but relatively approximate: We estimate $T_{c}$ by analyzing the behavior of the specific heat $C_{V}$ using three different methods for cross-validation. The second one involves estimating $T_{c}$ by analyzing the magnetization $M$ and magnetic susceptibility $\chi$. The third one offers greater accuracy and involves using graph convolutional networks (GCN) to refine our estimation of $T_{c}$.
@@
Our first method to estimate the specific heat utilizes the formula $C_{V}=\left(\frac{\partial E}{\partial T}\right)_{V}$. Given a stable spin configuration, the total energy $E$ of the system is calculated via Eq.(1). We then average the total energy over 10 stable spin configuration at each temperature, with a temperature interval $0.01J$. The numerical results are presented in the top panel of Fig.5. To estimate the specific heat, we use the spine functions to fit the $E$ versus $T$ curve, resulting in a function $E(T)$. We subsequently compute the derivative of this function to estimate $C_{V}$ and plot $C_{V}$ vs $T$ in the middle panel of Fig.5. As depicted in the figure, $C_{V}$ exhibits a singular behavior at $T/J\approx 2.29$, which can be considered as an estimate of $T_{c}$. Interestingly, this value is very close to the critical temperature of the two-dimensional Ising model on a square lattice, which is $T^{\Box}_{c}/J=2.269$ [31]. To quantify the uncertainty in this estimation of $T_{c}$, we apply a nonparametric Bootstrap resampling approach [32, 33], which has recently been applied to the Ising model [34]. Further details of the method are provided in Appendix A. The bootstrap iteration is performed $B=1000$ times, yielding a more precise estimate of $T_{c}$, namely $T_{c}=2.279\pm 0.008J$.
The second method for estimating $C_{V}$ is based on the the fluctuation-dissipation formula for specific heat:
$C_{V}=\frac{1}{T^{2}}\frac{\langle E^{2}\rangle-\langle E\rangle^{2}}{N},$ (3)
where the Boltzmann constant is set to 1 in natural units. Here, the energy $E$ is obtained using the previous method, and its absolute value is quite large. Therefore, when calculating $C_{V}$ by using Eq.(3), the squared terms become even larger, which can significantly amplify the oscillations of $E$. However, the results from the previous method may be influenced by the curve fitting technique. In contrast, this method does not require calculating the slope of the fitted curve, thereby avoiding errors caused by different fitting approaches. In the bottom panel of Fig.5, we present our calculations based on Eq.(3), where each data point is averaged over 10 stable spin patterns too. As expected, the value of $C_{V}$ exhibits significant oscillations as the temperature changes. However, there is also a distinct singularity around $T\approx 2.28J$. We further attempt to fit the curve using a Gaussian function and observe that the peak is indeed located around $2.28J$. To improve the accuracy of our results, we apply the Bootstrap error analysis, which refines the estimation of $T_{c}$ to $2.335\pm 0.292J$. Although this method introduces a larger uncertainty, the confidence interval of the results still overlaps with that obtained from the previous method.
Figure 6: (Top panel) Entropy as a function of temperature for the spherical ferromagnetic Ising model with $N=1000$. (Bottom panel) Specific heat as a function of temperature, calculated using $C_{V}=T\left(\frac{\partial S}{\partial T}\right)_{V}$. The inset shows the logarithmic fitting of data near $T_{c}$
[…]
Figure 7: (Top panel) Magnetization as a function of temperature for the spherical ferromagnetic Ising model with $N=1000$. (Bottom panel) Magnetic susceptibility as a function of temperature, calculated using $\chi=\frac{\langle M^{2}\rangle-\langle M\rangle^{2}}{NT}$.
The phase transition temperature $T_{c}$ can also be estimated by analyzing the variation of the magnetization $M$ and magnetic susceptibility $\chi$ with temperature. Given a stable spin configuration, the magnetization $M$ and susceptibility $\chi$ of the system can be calculated by
$M=\sum_{i}s_{i},\qquad\chi=\frac{\langle M^{2}\rangle-\langle M\rangle^{2}}{NT}.$ (4)
Through methods similar to those used for analyzing the temperature dependence of energy and specific heat, we average the total magnetization and susceptibility over 10 stable spin configuration at each temperature, with a temperature interval $0.01J$. The numerical results are presented in Fig.7, where the top panel shows the variation of the magnetization with temperature, and the bottom panel shows the variation of the susceptibility with temperature. As depicted in the figure, the susceptibility $\chi$ exhibits a singular behavior at $T/J\approx 2.33$, which can be considered as an estimate of $T_{c}$. We also apply the Bootstrap error analysis to improve the accuracy of the results, which refines the estimation of $T_{c}$ to $2.316\pm 0.053J$, which is close to the estimates of $T_{c}$ obtained from the specific heat analysis.
III.3 Determining $T_{c}$ via the method of machine learning
@@
The second approach employs machine learning, which has proven to be a powerful tool in processing big data [37, 38, 26, 39, 40, 41, 42, 43]. Given that the Fibonacci lattice is not homogeneous and does not resemble a typical image-like structure, traditional CNNs can not be applied to this model [44, 45]. Instead, we utilize GCNs, which are capable of capturing spatial structural features by leveraging the connectivity relationships between nodes [46]. This makes GCNs particularly well-suited for processing topologically structured data like the Fibonacci lattice. Details of GCN implementation can be found in Appendix C.
@@
Figure 8: Classification confidences, $p_{\text{o}}$ and $p_{\text{d}}$, versus temperature for the 2D Ising model on a 20$\times$20 square lattice (Top panel) and 40$\times$40 square lattice (Bottom panel).
To determine the critical temperature $T_{c}$ using GCNs, we adopt a supervised learning approach inspired by prior studies [33], but tailored to our system. Unlike traditional methods that rely on training samples labeled by a classifier with prior knowledge of $T_{c}$, we take advantage of the clear separation of phases at extreme temperatures to construct our training set. Specifically, Monte Carlo (MC) spin configurations are generated at very low temperatures (e.g., $T/J=0.01$, well below the expected $T_{c}$) and very high temperatures (e.g., $T/J=8.0$, well above the expected $T_{c}$), where the system is unambiguously in the ordered and disordered phases, respectively. These configurations are labeled as “ordered” ($p_{\text{o}}=1$, $p_{\text{d}}=0$) and “disordered” ($p_{\text{o}}=0$, $p_{\text{d}}=1$) based on their temperature regimes, eliminating the need for an exact $T_{c}$ value a priori. Here $p_{\text{o}}$ and $p_{\text{d}}$ represent the classification confidences for the ordered and disordered phases, respectively. The trained GCN then interpolates across the temperature range $[0.01,8.0]J$, predicting classification confidences $p_{\text{o}}$ and $p_{\text{d}}$ for each temperature. The intersection point where $p_{\text{o}}=p_{\text{d}}$ is taken as the estimated $T_{c}$.